# Asymmetric Contrastive Multimodal Learning for Advancing Chemical Understanding

### v4

@@ G-1H NMR GIN 5 512 256
B.2.2 Finetuning on Molecular Property Prediction Tasks
The following configurations were applied to all finetuning tasks from MoleculeNet [54] and TDC [55] accessed on July 15, 2023. For each task, we utilized a 5-layer GIN [57] as the backbone graph encoder, exploring hidden dimensions of [64, 128, 300]. The GIN encoder was initialized with weights pre-trained using the ACML framework. We used a batch size of 32 and a maximal training epoch of 100. We use scaffold splitting [86] to divide the dataset into the training set, validation set, and test set with a ratio of 8:1:1. Scaffold splitting helps avoid data leakage in molecular datasets because it ensures that structurally similar molecules do not appear in both the training and test sets. All experiments were conducted on one Tesla V100 GPU. We used the Adam optimizer. Different learning rates from [1e-2, 1e-3, 1e-4, 1e-5] are explored and the best results are reported. Before training, we performed data cleaning to remove certain molecules that failed to pass the sanitizing process in the RDKit or contained abnormal valence of a certain atom, as suggested in [87, 88, 44]. The detailed dataset statistics are summarized in Table 8.
@@ B.2.2 Finetuning on Molecular Property Prediction Tasks
B.2.3 Discussion of Training Efficiency
@@
Table 10: Resource Utilization and Time per Epoch Comparison
@@ G-1H NMR 1 V100, 32G $\sim$ 18K $\sim$ 12 sec
Appendix C Competing Interests
All authors have no competing interests to declare.
Appendix D Declaration of Generative AI and AI-assisted Technologies in The Writing Process
During the preparation of this work, the authors used ChatGPT and Grammarly to grammar check and improve writing. After using this tool/service, the authors reviewed and edited the content as needed and took full responsibility for the content of the publication.
Appendix E Data Availability
The pre-training data from NP-MRD and MoNA, and the zero-shot PubChem dataset are publicly available online, well-organized, and can be directly downloaded. RDKit can assist in parsing the dataset into individual samples.
Appendix F Code Availability
The code accompanying this work is available Github: https://github.com/GainGod-Xu/ACMLProject. The supplementary materials S1 and S2 will be available when this work is formal published.

### v5

# bibliography

@@ G-1H NMR GIN 5 512 256
4.8.2 Finetuning on Molecular Property Prediction Tasks
The following configurations were applied to all finetuning tasks from MoleculeNet [wu2018moleculenet] and TDC [huang2022artificial] accessed on July 15, 2023. For each task, we utilized a 5-layer GIN [xu2018how] as the backbone graph encoder, exploring hidden dimensions of [64, 128, 300]. The GIN encoder was initialized with weights pre-trained using the ACML framework. We used a batch size of 32 and a maximal training epoch of 100. We use scaffold splitting [ramsundar2019deep] to divide the dataset into the training set, validation set, and test set with a ratio of 8:1:1. Scaffold splitting helps avoid data leakage in molecular datasets because it ensures that structurally similar molecules do not appear in both the training and test sets. All experiments were conducted on one Tesla V100 GPU. We used the Adam optimizer. Different learning rates from [1e-2, 1e-3, 1e-4, 1e-5] are explored and the best results are reported. Before training, we performed data cleaning to remove certain molecules that failed to pass the sanitizing process in the RDKit or contained abnormal valence of a certain atom, as suggested in [ijcai2021, lim2021predicting, wang2023motif]. The detailed dataset statistics are summarized in Table 6.
@@ 4.8.2 Finetuning on Molecular Property Prediction Tasks
4.8.3 Discussion of Training Efficiency
@@
Table 8: Resource Utilization and Time per Epoch Comparison.
@@ G-1H NMR 1 V100, 32G $\sim$ 18K $\sim$ 12 sec
Supplemental Figures
Figure S1: 13C NMR Isomer discrimination task demonstrations. The correct pairs of 13C NMR and molecule are presented horizontally, linked by solid arrows. Out of the 4 challenging isomer pairs, our proposed model correctly identifies 3 isomers.
Figure S2: 1H NMR Isomer discrimination demonstrations. The correct pairs of 1H NMR and molecule are presented horizontally, linked by solid arrows. Out of the 4 challenging isomer pairs, our proposed model correctly identifies 3 isomers.
Figure S3: Graph embedding visualization via PCA. For each property, one representative figure is shown using the ACML instantiation that achieves the highest PCC score.